# FoAM: Foresight-Augmented Multi-Task Imitation Policy for Robotic Manipulation

### v1

@@ IV Experiments
Our experiments focus on investigating the following questions:
@@
    How does FoAM perform, and how does it outperform the baseline across multiple tasks?
@@
    What advantages does FoAM provide compared to a single goal-conditioned policy?
@@
    How effectively does FoAM handle unseen tasks without any data augmentation?
@@
    When FoAM and the VLM are used jointly during inference, how does FoAM perform differently when guided by an imagined goal image versus a real one?
@@
    A robust imitation learning policy is essential for achieving sustainable scalability. How well does FoAM respond to external disturbance?
@@
TABLE I: Performance of multi-task policies in our benchmark. In the table, the first column lists the names of the policies included in the evaluation, and the second column provides the model size of each policy. The subsequent six columns report the average success rates (See project homepage for the evaluation metrics of each task) of the respective policies across different task categories: dual-arm tasks, color block-based tasks, cabinet-based tasks, locker-based tasks, other tasks (including “Put something in the pan” and “Insert something in the basket”), and unseen tasks.
Figure 3: Snapshots of each scenario in the FoAM benchmark. The middle snapshot provides an overview of the dual-arm robot we developed in MuJoCo [55]. The tasks in the benchmark are categorized into five groups for performance analysis: pink for dual-arm tasks, yellow for cabinet-based tasks, green for color-block-based tasks, orange for locker-based tasks, and gray for other tasks. The objects in these scenarios are sourced from [56, 57, 58, 59]. The FoAM benchmark offers high-degree-of-freedom simulation suites. Tutorials for creating custom environments are available on the project homepage.
Figure 4: Snapshots of the real-world multi-task environment are captured from static externally mounted Orbbec Femto Bolt camera. We use a UFACTORY xArm 7 robot, a parallel-jaw gripper, a static externally mounted Orbbec Femto Bolt camera, and a wrist mounted Intel Realsense D435 camera, to evaluate multi-task policies across 4 real scenarios, comprising a total of 14 tasks. The tasks include: picking test tubes from the rack, placing fruits into the green bowl, inserting a test tube into holes (each with four subtasks), and placing the bitter melon on locker layers (with two subtasks). For videos and more details, please refer to the project homepage.
@@
TABLE II: The success rate of multi-task polices in real-world scenarios. Scenario I corresponds to the task: “Pick the first (second, third, forth) test tube from the rack.” Scenario II corresponds to: “Insert the test tube into the first (second, third, forth) hole.” Scenario III involves: “Put the eggplant (bitter melon, peach, tomato) in the green bowl.” Scenario IV refers to the task: “Place the bitter melon on the bottom (middle) layer of the locker.”
@@ IV-A Data Collection
FoAM Benchmark: We developed a simulated dual-arm robotic system, with each arm possessing 6 degrees of freedom (DoF) and a 1-dof parallel-jaw gripper, closely replicating a commonly used UR3e robot. This system was implemented with the MuJoCo physics simulation engine [55], and we have open-sourced 10 distinct multi-task suites involving this robot. A total of 86 simulation tasks were designed, encompassing a broad range of practical skills, such as picking, moving, pushing, placing, sliding, inserting, opening, closing, and transferring [25]. Figure 3 provides an overview of snapshots from each multi-task environment along with their corresponding scenario names. Each multi-task simulation environment includes varying numbers of subtasks. For example, the Open Cabinet Drawer scenario consists of three subtasks, with a general task prompt “Open the cabinet bottom drawer”, where “bottom” can be replaced with middle or top. The subtasks in the scenarios Transfer Color Blocks and Dual Arm: Put Stuff to the Cabinet Bottom Drawer are dual-arm tasks, while the remaining suites involve single-arm tasks. The FoAM benchmark is a high-degree-of-freedom simulation data generator, enabling users to customize textures, colors, and even action trajectories for tasks. This tool facilitates the rapid generation of high-quality simulation data tailored to user requirements. We expect that FoAM contributes to the development of multi-task policies in complex environments.
@@
Simulation Dataset: We generate 50 expert demonstrations for each task to evaluate the performance of FoAM versus other multi-task polices. Before recording each demonstration, objects in the scene are randomly initialized within a specified range. The dataset is recorded at a frequency of 50 Hz, capturing the robot’s proprioceptive data, action sequences, and visual observations. To simulate a real-world scenario where the agent interacts with users, visual observations are captured solely through a head-mounted camera with a resolution of $480\times 640$ pixels. For the single-arm tasks, even though one of the robotic arms remained inactive, we still recorded its action and proprioceptive data. As a result, the controlled action space $n$ of the dataset is unified to $14$, allowing us to accommodate all tasks within a single IL policy.
@@
Real World Dataset: Our robot system is composed of a UFACTORY xArm 7 robotic arm, a parallel-jaw gripper, a static externally mounted Orbbec Femto Bolt camera, and a wrist mounted Intel Realsense D435 camera. To evaluate the performance of FoAM in real-world applications, we designed 14 tasks across four multi-task environments. The 4 real-world scenarios are illustrated in Figure 4. The dataset was collected using a custom-built, low-cost teleoperation platform inspired by Gello [60]. The leading arm, which a teleoperator directly controls, includes eight Dynamixel XL330-M288-T servos and custom 3D-printed connectors (see project homepage for details). For each task, we collected 50 expert demonstrations, with the objects randomly placed on the table before data collection. The randomization was constrained within a rectangular area measuring approximately 50$\times$60 cm. The final dataset comprises RGB data from two cameras with a resolution of 480$\times$640, along with joint states from both the leading and following arms, recorded at a frequency of 30 Hz.
@@ IV-B Experiment Results
We compare our approach against state-of-the-art open-source multi-task policies, including Multi-task Action Chunking Transformer (MT-ACT) [9] and BAKU with a deterministic policy head [10], both of which utilize only language prompts. Additionally, we use ACT with goal images (Gimg-ACT) as a baseline guided solely by the goal image. For multi-modal prompting, ACT with both language prompts and goal images serves as a baseline policy, which can also function as an ablation policy to assess the effectiveness of the FA module (FoMA w/o FA).
@@
To facilitate statistical analysis, all scenarios in the FoAM benchmark were categorized into five distinct task groups. We conducted 50 test trials for each subtask, and the average success rates of the different strategies across these task categories are presented in Table I.
Compared to all the policies evaluated, FoAM achieved the highest success rate across all five task categories. Notably, in the dual-arm task, FoAM outperformed the second-best policy by 41% in success rate, with varying degrees of improvement observed in the other task categories as well. To further assess the generalization capabilities of these policies, we designed 4 unseen tasks by modifying the Scenario Pick Color Blocks: green blocks were changed to purple, and blue blocks to black. The language-based MT policies were unable to complete these unseen tasks, while policies incorporating goal images demonstrated varying levels of generalization. We think the reason is language-based policies rely on text embeddings to activate conditional responses, classifying tasks based on broad categories. This approach leaves the agent struggling when faced with unseen task prompts, as it lacks the flexibility to adapt to unseen tasks. In contrast, policies guided by goal images employ a more fine-grained classification approach, capturing detailed information from individual demonstrations. This enables the policy to focus more precisely on the differences between visual observations and goals, enhancing its ability to tackle unseen tasks. However, using a single goal image also presents limitations. In block-based scenarios, such as Transfer Color Blocks and Pick Color Blocks, where the initial and goal images are identical, this similarity can trigger the incorrect activation of the conditional policy. As a result, the Gimg-ACT performs significantly worse in these tasks compared to other polices. FoAM, by fusing both language and image inputs, provides stable activation conditions for each task while retaining the generalization advantages offered by goal images. When enhanced with FA, FoAM’s task performance is significantly improved.
Based on the performance of MT polices in FoAM benchmark, we strategically selected MT-ACT, Gimg-ACT, FoAM, and FoAM w/o FA for real-world deployment. In the real-world experiments, for each scenario, we randomly initialized ten different locations and sequentially executed the tasks associated with each scenario. The performance of these multi-task policies in real-world scenarios are summarized in Table II.
Each policy experienced a notable performance decline when deployed in real-world environments. We attribute this to the inherent randomness of human demonstrations and the variability present in real-world conditions, both of which complicate the processes of learning and inference. Furthermore, unlike the experimental environments used in prior works [20, 9], our scenarios closely mirror real-world robotic applications. Specifically, we utilized only two perspectives—head and wrist—and the manipulated objects were randomly initialized within a large workspace, increasing the challenge for the agent to complete tasks effectively. Scenarios I and II demand higher precision from the policy, as the test tubes and racket holes are closely located, making it sensitive to robot misoperations. In Scenario III, the 4 fruits and the green bowl were placed randomly, resulting in complex visual observations. In contrast, Scenario IV was relatively simpler for robot execution, involving larger manipulated objects and a well-defined goal space. However, this scenario also introduced goal image ambiguity, leading it difficult for Gimg-ACT to reliably activate specific tasks. Notably, FoAM demonstrated the best performance across all 4 scenarios. To assess policy performance on unseen tasks, we replaced the eggplant in Scenario III with a carambola. MT-ACT struggled achieving any success in the new tasks without data augmentation; in contrast, the other three goal-conditioned policies exhibited varying degrees of generalization, and FoAM achieves the highest success rate.
@@ IV-C VLM-FoAM Joint Inference
To improve the agent’s autonomy in acquiring the goal image, we conducted a joint inference experiment in Scenario III. Two policies were employed using data exclusively from Scenario III: FoAM, trained with the last frame of demonstration and evaluated with real goal images, and VLM-FoAM, trained and evaluated with goal images generated by VLM. The results of this experiment are shown in Table III.
Due to their shapes, Peach and Tomato are difficult for the robot to grasp and are prone to rolling, often leading to task failure. In contrast, Bitter Melon and Eggplant are more easily grasped. Our experiments showed that VLM-FoAM demonstrated more robust performance. We attribute this to the deep semantic information retained in the images generated by VLM, which helps prevent the model from overfitting when working with small datasets. Furthermore, the goal images generated by VLM maintain a consistent overall style. This style uniformity ensures that goal images generated at different times share similar features, enhancing the robot’s ability to adapt to dynamic real-world conditions, thereby improving task activation reliability. Additionally, with the introduction of VLM, the agent can autonomously and efficiently acquire the goal image, with a 480$\times$640 pixel goal image being obtained in an average of 4 seconds.
@@ V Conclusion and Future Work
In this work, we introduced FoAM, a novel multimodal goal-conditioned policy designed to enhance the performance of multi-task policies and address the limitations of single goal-conditioned approaches. Inspired by human behavior, FoAM improves agent performance by imitating expert actions while simultaneously considering the visual outcomes of those actions. In our published FoAM-benchmark and across real-world scenarios, FoAM achieved improvements of up to 41% in success rate compared with previous methods. However, FoAM exhibited certain limitations in real-world Scenarios I and II, which involve high precision requirements. To address this, we will explore to refine long-horizon tasks by generating fine-grained intermediate goal images to serve as guidance. By leveraging these intermediate visual states, we seek to reduce cumulative errors during operations and improve the agent’s execution accuracy.

### v2

@@ IV Experiments
Our experiments will answer the following questions:
@@
    How does multi-modal goal condition perform?
@@
    How does foresight augmentation perform?
@@
    Can FoAM generalize to unseen tasks with a few expert episodes and without data augmentation?
@@
    How does FoAM perform differently when guided by a real or generated goal image?
@@
    How well does FoAM respond to external disturbance?
@@
TABLE I: Performance of interested policies in FoAM benchmark. In the table, the first column lists the names of the policies included in the evaluation, and the second column provides the model size of each policy. The subsequent six columns report the average success rates of the respective policies across five task categories and four unseen tasks.
@@
TABLE II: Performance of interested policies in real-world scenarios. Scenario I corresponds to the task: “Pick the first (second, third, forth) test tube from the rack.” Scenario II corresponds to: “Insert the test tube into the first (second, third, forth) hole.” Scenario III involves: “Put the eggplant (bitter melon, peach, tomato) in the green bowl.” Scenario IV refers to the task: “Place the bitter melon on the bottom (middle) locker layer.”
@@ IV-A Data Collection
FoAM Benchmark. We developed a dual-arm system in MuJoCo[20], a popular physics simulation engine, with 6 degrees of freedom (DoF) each arm and a 1-DOF parallel-jaw gripper. This system is the first open-source simulated dual-arm system with rich tasks, manipulation scenarios, and the same physics as the real UR3e robot. It will serve as a valuable tool for developing zero-shot Sim2Real robotic reinforcement learning policies, such as DextrAH-G[21]. We have designed 10 distinct multi-task suites in the system. A total of 86 simulation tasks were involved, encompassing a broad range of practical skills, such as picking, moving, pushing, placing, sliding, inserting, opening, closing, and transferring [27]. Figure 3 provides an overview of snapshots from each multi-task scenario along with their corresponding names. Each scenario includes varying numbers of subtasks. For example, the Open Cabinet Drawer scenario consists of three subtasks, with a general task instruction “Open the cabinet bottom drawer”, where “bottom” can be replaced with “middle” or “top”. The subtasks in the scenarios Transfer Color Blocks and Dual Arm: Put Stuff to the Cabinet Bottom Drawer are dual-arm tasks, while the remaining suites involve single-arm tasks. The FoAM benchmark is a high-degree-of-freedom simulation data generator, enabling users to customize textures, colors, and even trajectories. This tool facilitates the rapid generation of high-quality simulation data tailored to user requirements by running scripts. The FoAM benchmark will enrich the existing simulation benchmark libraries, such as RLBench[57] and LIBERO [58]. We expect that FoAM will contribute to the development of multi-task policies in complex scenarios.
@@
Simulation Dataset. We generate 50 episodes for each task. Before recording each demonstration, objects in the scene are randomly initialized within a specified range. The dataset is recorded at a frequency of 50 Hz, capturing the robot’s proprioceptive data, action sequences, and visual observations. The visual observations are captured solely through a head-mounted camera with a resolution of $480\times 640$ pixels. For the single-arm tasks, even though one of the robotic arms remained inactive, we still recorded its action and proprioceptive data. As a result, the controlled action space $n$ of the dataset is unified to $14$, allowing us to accommodate all tasks within a single MTIL policy.
@@
Real World Dataset. Our robot system is composed of a UFACTORY xArm 7 robotic arm, a parallel-jaw gripper, a static externally mounted Orbbec Femto Bolt camera, and a wrist-mounted Intel Realsense D435 camera. To evaluate the performance of FoAM in the real world, we designed 14 tasks across four multi-task scenarios in the real world. The snapshots of four scenarios are illustrated in Figure 4. The dataset was collected using a custom-built, low-cost teleoperation platform inspired by Gello [59]. We collected 50 episodes for each task, with the objects randomly placed on the table before data collection. The randomization was constrained within a rectangular area measuring approximately 50$\times$60 cm. The final dataset comprises RGB data from two cameras with a resolution of 480$\times$640 each, along with joint states from both the leading and following arms, recorded at a frequency of 30 Hz.
Figure 3: Snapshots of each scenario in the FoAM benchmark. The middle snapshot provides an overview of the simulated dual-arm system we developed in MuJoCo [20]. The tasks in the benchmark are divided into five categories for evaluation: pink for dual-arm tasks, yellow for cabinet-based tasks, green for color-block-based tasks, orange for locker-based tasks, and gray for other tasks. The objects in these scenarios are sourced from [60, 61, 62, 63]. The FoAM benchmark offers high-degree-of-freedom simulation suites. Tutorials for creating custom environments are available on the project homepage.
@@ IV-B Experiment Results
We compare FoAM against state-of-the-art open-source MTIL policies, including Multi-task Action Chunking Transformer (MT-ACT) [9] and BAKU with a deterministic policy head [10], both of which utilize only task instruction as the goal condition. We evaluate ACT with goal images (Gimg-ACT) as a baseline guided solely by the goal image. The baseline of ACT with both task instruction and goal image serves as the multi-modal goal-conditioned policy, which can also function as an ablation experiment to assess the effectiveness of the foresight augmentation (FoMA w/o FA).
@@
All scenarios in the FoAM benchmark were categorized into five distinct task categories and four unseen tasks. We conducted 50 test trials for each task, and the average success rates of the different policies across these categories are presented in Table I. Compared to all the policies evaluated, FoAM achieved the highest success rate across all task categories. Notably, in dual-arm tasks, FoAM outperformed the second-best policy (Gimg-ACT) by 41% in success rate, with varying degrees of improvement observed in the other task categories as well.
Figure 4: Snapshots of the real-world multi-task environment are captured from static externally mounted Orbbec Femto Bolt camera. The tasks include Pick test tubes from the rack, Put fruits into the green bowl, Insert the test tube into holes (four subtasks each), and Place the bitter melon on locker layers (two subtasks). Please refer to the project homepage for videos and more details.
Advantages of multi-modal goal condition. The multi-modal goal condition enables the agent to perform unseen tasks with limited expert demonstrations and without data augmentation while solving the scenario ambiguity involved in the goal-image-based uni-modal policy. To evaluate the generalization capabilities of interested policies, we designed four unseen tasks by modifying the Scenario Pick Color Blocks: green blocks were changed to purple, and blue blocks to black. The language-based uni-modal policies (MT-ACT, BAKU) could not complete these unseen tasks, while policies involving goal image (Gimg-ACT, FoAM, FoAM w/o FA) demonstrated varying levels of generalization. We attribute this to the language-based policies that rely on text embeddings to conduct conditional responses, executing tasks based on the broad sub-tasks level. These policies leave the agent struggling when faced with unseen task instructions. In contrast, image-based uni-modal policies take actions based on pixel information, aligning goal information with a single episode, which would provide a more fine-grained representation on the episode level. This enables the policy to focus on precise pixel differences between visual observations and goals, allowing it to tackle unseen tasks.
The goal-image-based uni-modal policy suffers from scenario ambiguity, making it difficult to align tasks with intentions. For example, the scenarios in Block-based category, such as Transfer Color Blocks and Pick Color Blocks, have the same initial observation and goal images. The tasks, such as Transfer the right yellow block and Pick the right yellow block, have the same initial observation and goal image, which is the visual observation after the right yellow block disappears in the filed of view. Therefore, Gimg-ACT has the worst performance in Block-based tasks compared to other policies. The multi-modal goal condition of FoAM provides reliable execution conditions for each task while retaining the generalization ability offered by goal images.
Advantages of Foresight Augmentation. To discuss the benefits brought by Foresight Augmentation (FA), we designed the ablation item FoAM w/o FA. A comparative analysis of success rates across all five task categories reveals a significant performance enhancement in task execution after integrating the FA. We attribute this improvement to the policy’s ability to learn more expressive representations by aligning actions with their consequences, thereby facilitating more accurate task execution. According to Table I, our experiment results indicate that the ability to perform unseen tasks follows the order FoAM$>$Gimg-ACT$>$FoAM w/o FA, while language-based goal-conditioned policies (MT-ACT and BAKU) fail to generalize to unseen tasks. These results suggest that incorporating task instructions into the multi-modal goal condition influences the generalization capability conferred by the goal image. However, the integration of FA mitigates the negative impact of task instructions on generalization while enhancing the acquisition of more expressive representations, ultimately improving the policy’s overall generalization performance.
@@
Real-world experiment. Based on the performance of interested policies in FoAM benchmark, we strategically selected MT-ACT, Gimg-ACT, FoAM, and FoAM w/o FA for real-world deployment. In the real-world experiments, for each scenario, we randomly initialized ten different locations and sequentially executed the tasks associated with each scenario. The performance of these MTIL policies in real-world scenarios are summarized in Table II.
Each policy experienced a notable performance decline when deployed in real-world environments. We attribute this to the inherent randomness of expert demonstrations, distribution differences in multiple tasks, and the variability present in the dynamic real world, all of which complicate the processes of learning and inference. Furthermore, we utilized only two perspectives, head and wrist, and the manipulated objects were randomly initialized within a large workspace, increasing the challenge for the agent to learn skills effectively. Scenarios I and II demand higher accuracy from the actions, as the test tubes and racket holes are closely located, making them sensitive to robot misoperations. In Scenario III, the four fruits and the green bowl were placed randomly, resulting in complex visual observations. In contrast, Scenario IV was relatively simple for robot execution, involving just one large manipulated object and a well-defined goal space. However, Scenario IV introduced ambiguity, making it difficult for Gimg-ACT to reliably execute tasks consistent with intention.
Consistent with the simulation results, FoAM demonstrates superior performance across nearly all four real world scenarios. To evaluate policy performance on unseen tasks, we replaced the eggplant in Scenario III with a carambola. MT-ACT still struggled to achieve any success in the new tasks. The other three policies exhibited varying degrees of generalization, and FoAM achieved the highest success rate.
@@ IV-C VLM-FoAM Joint Inference
To improve the agent’s autonomy in acquiring the goal image, we conducted a joint inference experiment in Scenario III. Two policies were employed using data exclusively from Scenario III: FoAM, trained with the last frame of demonstration and evaluated with real goal images, and VLM-FoAM, trained and evaluated with goal images generated by VLM. The experiment results are shown in Table III.
Due to their shapes, Peach and Tomato are prone to rolling, which is difficult for the robot to grasp, leading to task failure. In contrast, Bitter Melon and Eggplant are more easily grasped. Although the experimental results have a small gap, we observed that VLM-FoAM exhibits more robust performance during the experiments. We attribute this to the deep semantic information retained in the images generated by VLM, which helps prevent the model from overfitting when working with small datasets. Furthermore, the goal images generated by VLM maintain a consistent overall style. This style uniformity ensures that goal images generated at different times share similar features, enhancing the robot’s ability to adapt to the dynamic real world, thereby improving task execution reliability. Additionally, with the introduction of VLM, the agent can autonomously and efficiently acquire the goal image, with a 480$\times$640 pixel goal image being obtained in an average of 4 seconds.
Although VLM brings benefits to robotic manipulation, we found some limitations during the experiment. When executing tasks in scenarios with complex workspaces and small manipulated objects, such as Scenario I and Scenario II, the generated goal images exhibited instability in semantic alignment. For example, when given the instruction Insert the test tube into the second hole, the generated image sometimes corresponded to an incorrect but visually similar placement, such as Inserting the test tube into the first/third hole. We attribute these discrepancies to the limited training data and the limitations of current VLMs in processing fine-grained pixel-level details. For instance, the test tube’s features constitute only a small portion of the overall visual input, posing a significant challenge for the model in accurately capturing and distinguishing such fine-grained spatial relationships.
@@ V Conclusion and Future Work
In this work, we introduced FoAM, a novel multi-modal goal-conditioned policy designed to enhance the performance of multi-task policies and address the limitations of uni-modal ones. Inspired by human action perception, FOAM improves the agent’s performance by mimicking expert actions while aligning actions with their consequences. In our published FoAM benchmark and across real-world scenarios, FoAM achieved improvements of up to 41% in success rate compared with previous policies. However, FoAM exhibited certain limitations in real-world Scenarios I and II, which involve high precision requirements. To address this, we will explore refining long-horizon tasks by generating fine-grained intermediate goal images to serve as consequence controllers. By leveraging them, we seek to reduce cumulative errors during manipulations and improve the agent’s execution accuracy.